OpenStudy (anonymous):

what is the missing integer -40/=5

OpenStudy (anonymous):

ask yourself: what can I multiply 5 by to find -40?

OpenStudy (anonymous):

alternatively you could just divide -40 by 5. your equation would look like this: -40/x=5. multiply both sides by x: -40=5x. then divide both sides by 5 to solve for x. x=-40/5

OpenStudy (goldphenoix):

It's -8. Think about it. \[\large \large \frac{ -40 }{ x } = 5\] What would the denonimator be... To make the equation better, you can rearrange the equation. Another way of writing this equation is, \[\large \large 5x = -40\] To solve for x, you want to divide both side by 5. So what is -40 divided by 5? Well, this is how you do it. Pretend that the negative sign is not there. What is 40 divided by 5? The answer to that question is 8. Now to see if 8 is a positive, you want to see if both number is either positive or negative. You have a positive 5, and a negative 40. This mean that 8 is a negative. The only way 8 can be a negative is if one of the number is positive, and the other number is negative. So your answer is -8. I hope this help!